OpenStudy (anonymous):

a building has a ramp to its front door. If the distance from the buidling to the end of the ramp is 20 feet and the height from the ground to the front doors is 4 ffet how long is the ramp

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) use the Pythagorean theorem a^2 + b^2 = c^2 in this case, a would equal 20 and b would equal 4 \[c = \sqrt{(20)^2+(4)^2} =~?\]